A bullet leaves a rifle with a muzzle velocity of 521 m/s. While accelerating through the barrel of the rifle, the bullet moves a distance of 0.840 m. Calculate the acceleration of the bullet.
1
Expert's answer
2020-12-17T09:17:00-0500

We can find the acceleration of the bullet from the kinematic equation:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{(521\ \dfrac{m}{s})^2-(0\ \dfrac{m}{s})^2}{2\cdot 0.84\ m}=1.62\cdot10^5\ \dfrac{m}{s^2}."

Answer:

"a=1.62\cdot10^5\ \dfrac{m}{s^2}."